Nicoley269:

How can I solve 3/1+root 5 ?

Hero:

@Nicoley269 would you mind taking a screenshot of the problem you're working on and upload it here?

Nnesha:

\[\frac{ 3 }{ 1+\sqrt{5} }\]

gabethebabe:

^what u posted is an expression is the question a rationalization question?

Nnesha:

multiply numerator and denominator by the conjugate of 1+sqrt{5}

Nicoley269:

What @Nnesha wrote is the the correct problem.

Nicoley269:

I don’t know how

Hero:

If that's the case, multiply top and bottom by \(1 - \sqrt{5}\)

gabethebabe:

do you understand the purpose of that step

Nicoley269:

I got up to\[3-3\sqrt{5} /1+\sqrt{5} *1-\sqrt{5}\]

Hero:

Great! So far you have \(\dfrac{3 - 3\sqrt{5}}{(1 + \sqrt{5})(1 - \sqrt{5})}\) All you have to do is just multiply the binomials in the denominator to do that you use the distributive property. Hopefully you're familiar with it: \(a(b + c) = ab + ac\) In this case, let \(a\) represent \(1 + \sqrt{5}\) \(b\) represent \(1\) \(c\) represent \(-\sqrt{5}\)

Hero:

I can show you an example of how to multiply two binomials if you like

Nicoley269:

I I don’t get it. Would the denominator be 2?

Nicoley269:

yes please

Hero:

Here's an example: \(\color\green{a}{(\color\red{b} + \color\purple{c})} = \color\green{a}\color\red{b} + \color\green{a}\color\purple{c}\) \( (\color\green{2 + \sqrt{6}})(\color\red{2} - \color\purple{\sqrt{6}}) \\= \color\red{2}\color\green{(2 + \sqrt{6})} \color\purple{- \sqrt{6}}\color\green{(2 + \sqrt{6})} \\=4 + 2\sqrt{6} - 2\sqrt{6} - 6 \\=4 - 6 \\=-2 \)

Hero:

Hope that's not too confusing for you

Nicoley269:

oh. so i was correct. 2

Hero:

That was just an example. I'd like to see you try multiplying the denominator on your own and posting all your steps here.

Nicoley269:

\[1(1-\sqrt{5})-\sqrt{5}(1+\sqrt{5})\]

Hero:

Mistakes in red: \(1(1\color\red{-}\sqrt{5})-\sqrt{5}(1+\sqrt{5})\)

Nicoley269:

i did minus.

Hero:

The binomials in parentheses must be the same in order to multiply correctly

Nicoley269:

ok...

Nicoley269:

i got 2

Hero:

Please show your work. All your steps.

Nicoley269:

nvm. its -4

Nicoley269:

 actually its 6

Nicoley269:

oops

Hero:

6 is correct

Hero:



Nicoley269:

thank you!